Circular No. 7440

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V4641 SAGITTARII J. A. Orosz, Utrecht University; E. Kuulkers, Space Research Organi-

zation Netherlands; M. van der Klis, University of Amsterdam; J. E. Mc-Clintock, Harvard-Smithsonian Center for Astrophysics; R. K. Jain and C. D. Bailyn, Yale University; and R. A. Remillard, Massachusetts Institute of Technology, report: “We have observed V4641 Sgr, the optical counterpart of the fast x-ray transient and superluminal radio source SAX J1819.3–2525 (IAUC 7119, 7253, 7254, 7256, 7265, 7271, 7277), on June 2–6 using the first 8.2-m telescope (+ FORS1) at the European Southern Observatory, Paranal. A preliminary analysis of the spectra indicates that the optical spectrum resembles that of an A2V star, with no evidence of emission lines from an accretion disk. We find an orbital period of 2.866 ± 0.043 days, a velocity semiamplitude of 204.7 ± 3.7 km/s, and an epoch of maximum velocity of HJD 2451701.658 ± 0.012. The resulting minimum mass of the compact object is 2.55 ± 0.14 M_{⊙}. Goranskij (1990, IBVS 3464) published the light curve of V4641 Sgr obtained from 345 plates taken with the Crimean 0.40-m astrograph between ∼ 1960 and 1990. Goranskij recently reanalyzed these data and made them available via the World Wide

Web (see http://www.kusastro.kyoto-u.ac.jp/vsnet/Mail/obs26000/

msg00925.html). We computed a Lomb-Scargle periodogram (cf. Scargle 1982, Ap.J. 263, 835, and cited references) of the data posted by Goranskij, excluding uncertain measurements and data from 1978 June, when the source apparently had another flare. There is significant power at a period of 1.40865 ± 0.000015 days. The data folded on twice this period, 2.8173 ± 0.00013 days, resembles an ellipsoidal light curve with two maxima of roughly equal height and two minima of unequal depth per orbital cycle. The epoch of minimum brightness is HJD 2447707.454 (IBVS 3464). The photometric phase at the time of maximum velocity is 0.741 ± 0.066, consistent with the expected value of 0.75 for an ellipsoidal light curve. The radius of the Roche-lobe-filling A2 star depends essentially only on its assumed mass. For an A2 star near the main sequence, M_{2} ∼ 2.3 M_{⊙} and the resulting radius is then ∼ 5.1 R_{⊙}. The implied distance is ∼ 6100 pc, assuming an effective temperature of 9000 K, an apparent magnitude of V = 14, and a color excess of E(B −V) = 0.24 (IAUC 7276), where the visual extinction is assumed to be 3.1E(B − V). Our current radial velocity curve is somewhat sparsely sampled, so we encourage additional spectroscopic observations.”

2000 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green